OpenStudy (anonymous):

Write in lowest terms 6x^2/12-6x / 2x^2-8x/x+2

OpenStudy (anonymous):

\[\frac{ 6x^2-24 }{ 12-6x }\div \frac{ 2x^2-8x }{ x+2 }\]

OpenStudy (anonymous):

\[\frac{ 5 }{ 8 }\div \frac{ 15 }{ 16 }-\frac{ 3 }{ 10 }\]